A 300kg particle has a velocity of (3.00i - 4.00j) m/s.

a). Find it's x and y components of momentum.

b). Find the magnitude and direction of its momentum.


1
Expert's answer
2021-05-31T08:39:32-0400

a)


"p_x=300(3)=900\frac{kgm}{s}\\p_y=-300(4)=-1200\frac{kgm}{s}"


b)


"p=\sqrt{900^2+1200^2}=1500\frac{kgm}{s}\\\theta=-\arctan{\frac{4}{3}}=-53\degree"